3. The production function for an economy can be expressed as Y FKL), where Y is real GDP, K is the quantity of capital in the economy, and L is the quantity of labor in the economy If F( ) = 100 + 3K+9L, what is real GDP if the quantity of capital is 200 and the quantity of labor is 500? b. What is/are the endogenous variable(s) in this model? c. What is/are the exogenous variable(s) in this model?
Answer #1

Answer 3

a)

Y = F(K,L) = 100 + 3K + 9L

where Y = rea GDP, K = quantity of Capita and L = quantity of Labor. It is given that K = 200 and L = 500

So, Y = 100 + 3K + 9L = 100 + 3*200 + 9*500 = 5200

Hence, Real GDP = 5200

b)

Endogenous variables are the Variables which are determined within the model and not considered as Given. In this question only Real GDP(Y) is not considered as given determined in the model and hence Endogenous variable is Real GDP

c)

Exogenous variables are the Variables which are considered as Given and are constant and are independent of any variable in the model. In this question Both K(capital) and Labor(L) are independent and considered as given and hence Exogenous variables are K(Capital) and L(Labor)